OpenStudy (anonymous):

sin 4x = (4sin x cosx)(2cos^2x-1)

OpenStudy (anonymous):

is this a proof? if so start with left hand side \[\sin(4x)\] \[2\sin(2x)\cos(2x)\] \[2(2sinxcox)(2\cos^2(x) - 1)\] \[(4sinxcox)(2\cos^2x-1)\]